find nth term of series 4,11,22,37,56
1
Expert's answer
2013-03-05T07:48:04-0500
we see that
a1=4
a2=4+7
a3=4+7+11
a4=4+7+11+15
a5=4+7+11+15+19

so a_n=a_(n-1)+4n-1=a_(n-2)+4*(n-1)-1+4*n-1=....=a1+4*2-1+4*3-1+....+4*n-1=
=a1+4*( 2+3+4+...+n) -(n-1)=a1+4*(2+n)*(n-1)/2-(n-1)=a1+2*(n+2)*(n-1)-(n-1)= =a1+(n-1)(
2(n+2)-1)=a1+(n-1)*(2n+3)=4+(n-1)(2n+3)
Hence a_n=4+2n^2+n-3=2n^2+n+1